# GDGB: A Benchmark for Generative Dynamic Text-Attributed Graph Learning

### v1

@@ 6 Future Work
Generative Framework Optimization. Based on our proposed high-quality and text-rich GDGB datasets, our generative framework has achieved groundbreaking results in two novel DyTAG generation tasks, including TDGG and IDGG. While these results demonstrate significant progress, several critical aspects still warrant refinement. Future research is expected to focus on optimizing the generation pipeline, particularly in IDGG, by refining node generation strategies and node selection mechanisms. For example, adaptive node sampling modules could balance diversity and fidelity, while current techniques in DGNNs could be appropriately used to improve accuracy and efficiency in recalling and ranking candidate nodes. Based on GDGB, these advancements towards a stronger DyTAG generative model would lay a robust foundation for graph foundation models in generative tasks, addressing challenges in both structural fidelity and attribute richness.
Further Applications. Beyond technical improvements, DyTAG generation offers significant practical applications. In Section 5.3, we emphasize that the identification of generated future hub nodes (e.g., potential hitting products in e-commerce) highlights its value in predictive modeling at the node level. Additionally, at the edge and graph levels, DyTAGs can act as practical data augmentation tools for sparse graph datasets, generating synthetic graphs that preserve structural properties while enriching textual attributes. Hence, the utility of DyTAG generation could be extended to diverse domains such as e-commerce (simulating consumer-product interaction dynamics), social network analysis (forecasting misinformation spread), and urban planning (predicting infrastructure usage patterns). By bridging dynamic graph analysis and forward-looking textual attribute modeling, DyTAG generation has the potential to facilitate real-world applications of various industries.
@@ 7 Conclusion
In this work, we propose Generative DyTAG Benchmark (GDGB), which resolves existing limitations in datasets, task formulations, and evaluation metrics for DyTAG generation. GDGB comprises eight meticulously curated datasets with semantically rich node/edge attributes, enabling rigorous evaluation of DyTAG generation. By defining two novel tasks—TDGG (transductive generation with fixed node sets) and IDGG (inductive generation with new node generation)—we effectively model the dynamic expansion of real-world graphs. Additionally, we propose GAG-General, an LLM-based multi-agent framework that generalizes across diverse DyTAG structures and integrates multifaceted metrics for holistic evaluation in DyTAG generation. Experimental results demonstrate that GDGB establishes robust benchmarks for DyTAG generation, revealing the interplay between structural and textual features as a cornerstone for effective generative models. In summary, GDGB lays the groundwork for future research on advancing DyTAG generation.
Acknowledgments and Disclosure of Funding
The work was partially done at Gaoling School of Artificial Intelligence, Beijing Key Laboratory of Research on Large Models and Intelligent Governance, MOE Key Lab of Data Engineering and Knowledge Engineering, Engineering Research Center of Next-Generation Intelligent Search and Recommendation, MOE, and Pazhou Laboratory (Huangpu), Guangzhou, Guangdong 510555, China. This research was supported in part by National Natural Science Foundation of China (No. 92470128, No. U2241212), by Beijing Outstanding Young Scientist Program No.BJJWZYJH012019100020098, by the National Key Research and Development Plan of China (2023YFB4502305) and Ant Group Research Fund. We also wish to acknowledge the support provided by the fund for building world-class universities (disciplines) of Renmin University of China, by Engineering Research Center of Next-Generation Intelligent Search and Recommendation, Ministry of Education, by Intelligent Social Governance Interdisciplinary Platform, Major Innovation & Planning Interdisciplinary Platform for the “Double-First Class” Initiative, Public Policy and Decision-making Research Lab, and Public Computing Cloud, Renmin University of China.
Appendix A Dataset
A.1 Dataset Details
@@
Dianping [^2] http://www.dianping.com dianping1 ; dianping2 is a business review dataset derived from Dianping, a prominent platform for business recommendations (e.g., restaurants), spanning from July 3, 2009, to February 8, 2012. The dataset records Dianping users’ historical review statistics, frequently reviewed cities, and detailed business information, including store names, addresses, cuisine styles, average costs, and historical scores. User reviews of businesses include multi-dimensional ratings (e.g., flavor, environment, service) in addition to overall scores (ranging from 0 to 5) as edge labels. Reviews also contain rich textual content and supplementary evaluations for specific items such as recommended dishes or special services, providing extensive edge textual features. The dataset is structured as a bipartite DyTAG, where Dianping users and businesses are nodes with textual features, and an edge represents a user’s detailed textual review of a business at a given time.
@@
WikiLife wikilife is a dataset derived from Wikipedia’s official data dumps, recording historical records of notable individuals being physically present at specific locations. The temporal span ranges from 202 to 2024. The original dataset wikilife is extracted from English Wikipedia’s biography pages as person-time-location triplets. We extend this by crawling the first paragraphs of corresponding Wikipedia pages for persons and locations as extra textual features. Additionally, we utilize 24 life trajectory activity categories (e.g., birth/death, education, career) from wikilife as edge labels and map the original Wikipedia text of triplets to describe the specific activities of individuals at locations. The WikiLife dataset is structured as a bipartite DyTAG, where persons and locations are nodes with textual features, and an edge represents a person’s textual life trajectory activity in a location at a given time.
@@
WeiboTechsagraph is a dataset collected from the Weibo social platform, recording user interactions (comments and reposts). The temporal span ranges from December 29, 2023, to January 5, 2024. The original dataset sagraph focuses on advertising strategies for electric toothbrushes. We reprocess the raw data to extract temporal information and restructure it into the DyTAG format. The dataset includes user profiles (e.g., usernames, gender, regions, follower/followee counts, self-introductions) and interaction edges categorized as comment or repost, with edge texts comprising source post content and destination user comments/reposts. Named WeiboTech due to its focus on technology-related topics (e.g., electronics, automobiles), the dataset is a non-bipartite DyTAG, where Weibo users are nodes with textual features, and an edge represents user interactions at a given time.
WeiboDailysagraph is also a dataset collected from the Weibo social platform, recording user interactions (comments and reposts). Different from WeiboTech, the temporal span ranges from December 1, 2023, to December 31, 2023, and the original dataset sagraph focuses on advertising strategies for ABCReading. Thus, WeiboDaily spans a full month and targets daily life topics (e.g., lifestyle sharing), enabling continuous modeling analysis. Similarly, the dataset includes user profiles (e.g., usernames, gender, regions, follower/followee counts, self-introductions) and interaction edges categorized as comment or repost, with edge texts comprising source post content and destination user comments/reposts. The dataset is a non-bipartite DyTAG, where Weibo users are nodes with textual features, and an edge represents user interactions at a given time.
Cora is an extended version of the classic citation network dataset Cora cora , documenting academic paper citation relationships. The temporal span ranges from February 1, 1985, to September 1, 2024. The original Cora dataset records citation relationships and node categories cora . We expand the dataset by crawling first-order and second-order paper information via references, enriching the size of the citation network. Node textual features include paper titles, abstracts, authors, and citation counts. Edge texts are extracted from the exact sentence in the citing paper that references the cited paper. With regards to the edge labels, based on the section name of each citation in the citing paper, they are mapped to five categories: 1) Intro & Background, 2) Tech & Methodology, 3) Experiment & Conclusion, 4) Topic-specific, and 5) Others. The Cora dataset is a non-bipartite DyTAG, where papers are nodes with textual features, and an edge indicates a citation relationship at a given time.
A.2 Dataset Licenses
@@
A.3 Dataset Analysis
To provide a comprehensive overview of our GDGB datasets, we present the node degree distribution in Figure 4. Most datasets exhibit a long-tailed distribution, which aligns with the real-world graph growth patterns governed by power-law principles powerlaw . This characteristic reflects the uneven connectivity observed in natural networks, such as social networks or citation graphs, where a few nodes dominate high degrees while the majority remain sparsely connected.
Furthermore, Figure 5 illustrates the edge label distribution of the GDGB datasets. The results reveal that most datasets maintain a balanced label distribution, mitigating the risk of performance degradation in edge classification tasks caused by extreme class imbalances. However, a subset of datasets exhibits relatively skewed label distributions, intentionally introducing realistic challenges akin to practical applications (e.g., rare but critical edge types in life trajectory or movie collaboration scenarios). This dual distribution pattern ensures both diversity and practical relevance, enhancing the dataset’s utility for robust model evaluation.
@@ A.3 Dataset Analysis
Figure 5: Distribution of the number of edges for each label on GDGN datasets.
A.4 Comparison with Existing Datasets for Graph Generation.
In graph generation research, while significant progress has been made in molecular graph generation, non-molecular graph generation has received comparatively less attention DBLP:journals/access/FaezOBR21 . Consequently, existing models for non-molecular graph generation are often trained on synthetic datasets such as Grid Bigg , Community, and SBM graphrnn ; DBLP:journals/pami/GuoZ23 , or on static, topology-only graphs, such as Ego and Polblogs EDGE . Further, we observe a lack of focus on dynamic graph generation, despite the fact that real-world non-molecular graphs inherently exhibit temporal evolution in both topology and node attributes DBLP:journals/pami/GuoZ23 . Among the most commonly used continuous-time dynamic graph datasets, such as LastFM and MOOC DGGEN ; TIGGER ; DGB , both are derived from interaction networks DBLP:conf/nips/PoursafaeiHPR22 that inherently include textual attributes of users/items and timestamped interactions. However, these datasets suffer from critical limitations: MOOC provides only numerical features extracted from raw text, while LastFM discards attribute information entirely.
Although widely adopted, existing datasets face notable challenges in modeling the co-evolution of temporal, structural, and textual attributes in graphs. First, static graph datasets lack temporal annotations, making them incompatible with dynamic graph generation tasks. Even discrete-time dynamic graph datasets like Bitcoin-Alpha bitcoin ; TIGGER , which usually represent graphs as discrete daily snapshots, suffer from sparse temporal distributions of edges, complicating the fine-grained modeling of structural and temporal dynamics. Second, most dynamic graph datasets omit raw textual data, relying instead on preprocessed numeric or categorical features DGGEN , which limits the exploration of rich textual attributes in real-world scenarios. Finally, more recently, datasets such as the DyTAGs from DTGB DTGB , while providing edge-level text features, often lack sufficient and high-quality node-level textual information. For instance, six out of eight datasets of DTGB contain (source) node texts that are typically just identifiers like email addresses or username. This is a critical limitation, as real-world applications frequently require detailed node-wise text attributes to disambiguate entities beyond IDs. The statistics of the above mentioned datasets and our proposed datasets are shown in Table 11.
@@
WikiRevision 75,622/3,204 2,778,732 2 276,6153 Web Interaction Node & Edge
@@
A.5 Dataset Documentation
The resource links for GDGB datasets, source codes, and leaderboards are provided as follows.
Appendix B Details of Tasks and Metrics
B.1 Details of TDGG and IDGG on GAG-General

### v2

@@ 6 Future Work
Generative Framework Optimization. Leveraging our high-quality, text-rich GDGB datasets, the proposed generative framework achieves groundbreaking performance in the novel DyTAG generation tasks of TDGG and IDGG. Despite this progress, key aspects of the pipeline—especially in IDGG—require further refinement. Future work should focus on improving node generation strategies and selection mechanisms, such as incorporating adaptive node sampling to balance diversity and fidelity, and leveraging existing DGNN techniques to enhance the accuracy and efficiency of candidate node retrieval and ranking. These advancements, grounded in GDGB, will strengthen DyTAG models and provide a solid foundation for generative graph foundation models, addressing challenges in both structural fidelity and attribute richness.
Further Applications. Beyond model optimization, DyTAG generation holds substantial practical value. As shown in Section˜5.3, identifying generated future hub nodes—like potential hit products in e-commerce—demonstrates its utility in node-level predictive modeling. At the edge and graph levels, DyTAG generation serves as effective data augmentation tools for sparse graphs, producing synthetic graphs that preserve structure while enriching textual content. This capability extends applications to e-commerce (modeling generative recommendation), social networks (forecasting misinformation spread), and urban planning (predicting infrastructure usage). By integrating dynamic graph analysis with forward-looking textual modeling, DyTAG generation can drive impactful real-world solutions across industries.
@@ 7 Conclusion
In this work, we propose Generative DyTAG Benchmark (GDGB), which resolves existing limitations in datasets, task formulations, and evaluation metrics for DyTAG generation. GDGB comprises 8 meticulously curated datasets with semantically rich node/edge attributes, enabling rigorous evaluation of DyTAG generation. By defining two novel tasks—TDGG (transductive generation with fixed node sets) and IDGG (inductive generation with new node generation)—we effectively model the dynamic expansion of real-world graphs. Additionally, we propose GAG-General, an LLM-based multi-agent framework that generalizes across diverse DyTAG structures and integrates multifaceted metrics for holistic evaluation in DyTAG generation. Experimental results demonstrate that GDGB establishes robust benchmarks for DyTAG generation, revealing the interplay between structural and textual features as a cornerstone for effective generative models. In summary, GDGB lays the groundwork for future research on advancing DyTAG generation.
Acknowledgments
The work was partially done at Gaoling School of Artificial Intelligence, Beijing Key Laboratory of Research on Large Models and Intelligent Governance, Engineering Research Center of Next-Generation Intelligent Search and Recommendation, MOE, and Pazhou Laboratory (Huangpu), Guangzhou, Guangdong 510555, China. This research was supported in part by National Natural Science Foundation of China (No. 92470128, No. U2241212), by Beijing Outstanding Young Scientist Program No.BJJWZYJH012019100020098, by the National Key Research and Development Plan of China (2023YFB4502305) and Ant Group through CCF-Ant Research Fund. We also wish to acknowledge the support provided by the fund for building world-class universities (disciplines) of Renmin University of China, by Engineering Research Center of Next-Generation Intelligent Search and Recommendation, Ministry of Education, by Intelligent Social Governance Interdisciplinary Platform, Major Innovation & Planning Interdisciplinary Platform for the “Double-First Class” Initiative, Public Policy and Decision-making Research Lab, and Public Computing Cloud, Renmin University of China.
Appendix A Ethics Statement
Appendix B Reproducibility Statement
To ensure the reproducibility of our work, we have made the following efforts: Our proposed GDGB comprises eight carefully selected and rigorously processed DyTAG datasets, hosted through a sharing link (https://www.kaggle.com/datasets/gdgbdataset/gdgb-a-benchmark-for-generative-dytag-learning). The proposed GDGB benchmark and our generative framework GAG-General are fully described in this paper, including detailed dataset construction processes and model implementation specifications. All code and data preprocessing scripts are publicly available in our code repository (https://github.com/Lucas-PJ/GDGB-ALGO), which includes comprehensive instructions for reproducing the TDGG and IDGG experiments. The repository also provides evaluation scripts of our proposed multifaceted metrics to facilitate verification of the reported results. The GDGB website is https://gdgb-algo.github.io/. The website allows researchers to submit methods and track performance via leaderboards on TDGG and IDGG.
Appendix C Dataset
C.1 Dataset Details
@@
Dianping [^2] http://www.dianping.com (Zhang et al., 2013; 2014) is a business review dataset derived from Dianping, a prominent platform for business recommendations (e.g., restaurants), spanning from July 3, 2009, to February 8, 2012. The dataset records Dianping users’ historical review statistics, frequently reviewed cities, and detailed business information, including store names, addresses, cuisine styles, average costs, and historical scores. User reviews of businesses include multi-dimensional ratings (e.g., flavor, environment, service) in addition to overall scores (ranging from 0 to 5) as edge labels. Reviews also contain rich textual content and supplementary evaluations for specific items such as recommended dishes or special services, providing extensive edge textual features. The dataset is structured as a bipartite DyTAG, where Dianping users and businesses are nodes with textual features, and an edge represents a user’s detailed textual review of a business at a given time.
@@
WikiLife (Zhang et al., 2024c) is a dataset derived from Wikipedia’s official data dumps, recording historical records of notable individuals being physically present at specific locations. The temporal span ranges from 202 to 2024. The original dataset (Zhang et al., 2024c) is extracted from English Wikipedia’s biography pages as person-time-location triplets. We extend this by crawling the first paragraphs of corresponding Wikipedia pages for persons and locations as extra textual features. Additionally, we utilize 24 life trajectory activity categories (e.g., birth/death, education, career) from (Zhang et al., 2024c) as edge labels and map the original Wikipedia text of triplets to describe the specific activities of individuals at locations. The WikiLife dataset is structured as a bipartite DyTAG, where persons and locations are nodes with textual features, and an edge represents a person’s textual life trajectory activity in a location at a given time.
@@
WeiboTech(Zhang et al., 2024b) is a dataset collected from the Weibo social platform, recording user interactions (comments and reposts). The temporal span ranges from December 29, 2023, to January 5, 2024. The original dataset (Zhang et al., 2024b) focuses on advertising strategies for electric toothbrushes. We reprocess the raw data to extract temporal information and restructure it into the DyTAG format. The dataset includes user profiles (e.g., usernames, gender, regions, follower/followee counts, self-introductions) and interaction edges categorized as comment or repost, with edge texts comprising source post content and destination user comments/reposts. Named WeiboTech due to its focus on technology-related topics (e.g., electronics, automobiles), the dataset is a non-bipartite DyTAG, where Weibo users are nodes with textual features, and an edge represents user interactions at a given time.
WeiboDaily(Zhang et al., 2024b) is also a dataset collected from the Weibo social platform, recording user interactions (comments and reposts). Different from WeiboTech, the temporal span ranges from December 1, 2023, to December 31, 2023, and the original dataset (Zhang et al., 2024b) focuses on advertising strategies for ABCReading. Thus, WeiboDaily spans a full month and targets daily life topics (e.g., lifestyle sharing), enabling continuous modeling analysis. Similarly, the dataset includes user profiles (e.g., usernames, gender, regions, follower/followee counts, self-introductions) and interaction edges categorized as comment or repost, with edge texts comprising source post content and destination user comments/reposts. The dataset is a non-bipartite DyTAG, where Weibo users are nodes with textual features, and an edge represents user interactions at a given time.
Cora is an extended version of the classic citation network dataset Cora (Sen et al., 2008), documenting academic paper citation relationships. The temporal span ranges from February 1, 1985, to September 1, 2024. The original Cora dataset records citation relationships and node categories (Sen et al., 2008). We expand the dataset by crawling first-order and second-order paper information via references, enriching the size of the citation network. Node textual features include paper titles, abstracts, authors, and citation counts. Edge texts are extracted from the exact sentence in the citing paper that references the cited paper. With regards to the edge labels, based on the section name of each citation in the citing paper, they are mapped to five categories: 1) Intro & Background, 2) Tech & Methodology, 3) Experiment & Conclusion, 4) Topic-specific, and 5) Others. The Cora dataset is a non-bipartite DyTAG, where papers are nodes with textual features, and an edge indicates a citation relationship at a given time.
C.2 Dataset Licenses
@@
C.3 Dataset Analysis
To provide a comprehensive overview of our GDGB datasets, we present the node degree distribution in Figure˜3. Most datasets exhibit a long-tailed distribution, which aligns with the real-world graph growth patterns governed by power-law principles (Clauset et al., 2009). This characteristic reflects the uneven connectivity observed in natural networks, such as social networks or citation graphs, where a few nodes dominate high degrees while the majority remain sparsely connected.
Furthermore, Figure˜4 illustrates the edge label distribution of the GDGB datasets. The results reveal that most datasets maintain a balanced label distribution, mitigating the risk of performance degradation in edge classification tasks caused by extreme class imbalances. However, a subset of datasets exhibits relatively skewed label distributions, intentionally introducing realistic challenges akin to practical applications (e.g., rare but critical edge types in life trajectory or movie collaboration scenarios). This dual distribution pattern ensures both diversity and practical relevance, enhancing the dataset’s utility for robust model evaluation.
@@ C.3 Dataset Analysis
Figure 4: Distribution of the number of edges for each label on GDGB datasets.
C.4 Comparison with Existing Datasets for Graph Generation.
In graph generation research, while significant progress has been made in molecular graph generation, non-molecular graph generation has received comparatively less attention (Faez et al., 2021). Consequently, existing models for non-molecular graph generation are often trained on synthetic datasets such as Grid (Dai et al., 2020), Community, and SBM (You et al., 2018; Guo and Zhao, 2023), or on static, topology-only graphs, such as Ego and Polblogs (Chen et al., 2023). Further, we observe a lack of focus on dynamic graph generation, despite the fact that real-world non-molecular graphs inherently exhibit temporal evolution in both topology and node attributes (Guo and Zhao, 2023). Among the most commonly used continuous-time dynamic graph datasets, such as LastFM and MOOC (Hosseini et al., 2024; Gupta et al., 2022; Poursafaei et al., 2022b), both are derived from interaction networks (Poursafaei et al., 2022a) that inherently include textual attributes of users/items and timestamped interactions. However, these datasets suffer from critical limitations: MOOC provides only numerical features extracted from raw text, while LastFM discards attribute information entirely.
Although widely adopted, existing datasets face notable challenges in modeling the co-evolution of temporal, structural, and textual attributes in graphs. First, static graph datasets lack temporal annotations, making them incompatible with dynamic graph generation tasks. Even discrete-time dynamic graph datasets like Bitcoin-Alpha (Kumar et al., 2016; Gupta et al., 2022), which usually represent graphs as discrete daily snapshots, suffer from sparse temporal distributions of edges, complicating the fine-grained modeling of structural and temporal dynamics. Second, most dynamic graph datasets omit raw textual data, relying instead on preprocessed numeric or categorical features (Hosseini et al., 2024), which limits the exploration of rich textual attributes in real-world scenarios. Finally, more recently, datasets such as the DyTAGs from DTGB (Zhang et al., 2024a), while providing edge-level text features, often lack sufficient and high-quality node-level textual information. For instance, six out of eight datasets of DTGB contain (source) node texts that are typically just identifiers like email addresses or username. This is a critical limitation, as real-world applications frequently require detailed node-wise text attributes to disambiguate entities beyond IDs. The statistics of the above mentioned datasets and our proposed datasets are shown in Table˜9.
@@
WikiRevision 75,622/3,204 2,778,732 2 2,766,153 Web Interaction Node & Edge
@@
C.5 Comparison Between GDGB and DTGB datasets
We provide a comprehensive analysis and empirical comparison between our proposed GDGB and the existing DTGB. While both benchmarks contain dynamic text-attributed graphs, we demonstrate that GDGB is fundamentally distinct in construction philosophy, data quality, and utility for generative tasks.
Figure 5: Left: Average node text lengths on GDGB and DTGB datasets. In non-bipartite graphs, text lengths are averaged across all nodes. For bipartite graphs, averages are calculated for source nodes. Right: Average edge text lengths on each dataset.
Figure 7: Left: Comparison of Degree MMD w/ and w/o text on GDGB datasets and DTGB datasets in VRDAG. Right: Comparison of Spectra MMD w/ and w/o text on GDGB datasets and DTGB datasets in VRDAG.
Textual Information Richness and Functional Utility.
As illustrated in Figure˜5, we compare the average text length of node and edge attributes across datasets in GDGB and DTGB, demonstrating that GDGB datasets exhibit significantly richer textual content. However, text length alone does not guarantee utility in generative modeling. To assess the functional value of textual features, we evaluate two feature-aware dynamic graph generation models—VRDAG (Li et al., 2024) and DG-Gen (Hosseini et al., 2024)—on both benchmarks, with and without the incorporation of BERT-based textual embeddings. As shown in Figures˜7 and 7, textual features consistently improve generation quality on GDGB, as evidenced by reduced Degree and Spectra MMD scores. In contrast, on DTGB, the inclusion of textual features often degrades performance—particularly for VRDAG, which is highly dependent on node feature inputs. This suggests that the textual attributes in DTGB are semantically impoverished or noisy, and thus detrimental to generative tasks, whereas GDGB’s textual content is both meaningful and effectively leveraged by learning models.
Intrinsic Text Quality Evaluation.
To directly evaluate the linguistic quality of node and edge texts, we conduct intrinsic analysis using two complementary metrics: (i) Perplexity (PPL), where lower values indicate higher fluency and coherence with natural language distributions; and (ii) LLM-based human-like scoring (on a 1–5 scale) across five dimensions: Contextual Fidelity, Personality Depth, Dynamic Adaptability, Immersive Quality, and Content Richness. The aggregated results are summarized in Table˜10. GDGB exhibits significantly lower perplexity (average value less than 90), indicating that its textual content is more coherent and aligned with natural language patterns. Furthermore, LLM evaluations confirm the superior semantic quality of GDGB, with average scores exceeding 4.0 across both node and edge texts. In contrast, DTGB achieves notably lower scores for node texts (average: 3.03), particularly in datasets such as Enron, where node attributes often consist of identifiers (e.g., email addresses) lacking meaningful semantics.
Table 10: Intrinsic text quality comparison (averaged across datasets).
Metric DTGB GDGB
PPL of node texts $\downarrow$ 193.58 75.95
PPL of edge texts $\downarrow$ 181.82 88.17
LLM rating (node texts) $\uparrow$ 3.03 4.21
LLM rating (edge texts) $\uparrow$ 3.55 4.17
Generative Performance on GAG-General.
To further assess the suitability of each benchmark for DyTAG generation, we apply the newly proposed GAG-General framework to all datasets from DTGB and all datasets from GDGB, using GPT as the LLM backbone under both TDGG and IDGG tasks. All models are evaluated under identical configurations in Section˜E.4 to ensure fair comparison. The results under TDGG are presented in Table˜11, 12, and 13. While, the results under IDGG are presented in Table˜14, 15, and 16. The above results reveal consistent advantages for GDGB across structural, textual, and embedding fidelity metrics. For instance, in the graph embedding metric under TDGG (Table˜13), DTGB results predominantly fall within the 0.2–0.4 range, while GDGB achieves substantially greater similarity (0.4–0.7), indicating better preservation of global structural and semantic properties. An even more pronounced trend is observed under IDGG: none of the DyTAGs generated on DTGB datasets satisfy the power-law validity criterion (Table˜14), and both average textual quality scores (Table˜15) and graph embedding metrics (Table˜16) are generally lower than those achieved on GDGB datasets. This again highlights the limitations of DTGB’s textual content in effectively supporting generative DyTAG tasks. Although the integration of memory and reflection mechanisms in GAG-General yields improvements on DTGB, these gains are constrained by the underlying data limitations, underscoring the dataset’s inherent constraints for generative tasks.
Table 11: The results of graph structural quality under TDGG with GPT as the LLM backbone.
Benchmark Dataset Degree MMD $\downarrow$ Spectra MMD $\downarrow$ $D_{k}$ $\alpha$ Power-law Validity
DTGB Enron 0.331 0.266 0.125 1.982 $\times$
GDELT 0.252 0.326 0.174 1.829 $\times$
ICEWS1819 0.302 0.341 0.132 1.794 $\times$
Stack elec 0.228 0.345 0.133 1.822 $\times$
Stack ubuntu 0.235 0.359 0.130 1.901 $\times$
Googlemap CT 0.255 0.375 0.080 2.278
Amazon movies 0.404 0.350 0.182 1.636 $\times$
Yelp 0.213 0.224 0.121 2.033
GDGB Sephora 0.023 0.011 0.143 2.993
Dianping 0.055 0.328 0.041 2.234 $\checkmark$
WikiRevision 0.108 0.156 0.056 2.041 $\checkmark$
WikiLife 0.181 0.223 0.099 2.204
IMDB 0.278 0.316 0.135 1.720 $\times$
WeiboTech 0.243 0.297 0.030 2.011
WeiboDaily 0.247 0.493 0.048 1.845 $\times$
Cora 0.128 0.156 0.049 2.378
Table 12: The results on average textual quality scores under TDGG with GPT as the LLM backbone. M. and R. denote node memory and reflection mechanism, respectively.
Benchmark Dataset w/o M. w/ M. w/ M. & R.
DTGB Enron 3.65 3.72 3.83
GDELT 3.70 3.98 4.23
ICEWS1819 3.47 3.87 4.14
Stack elec 3.68 3.70 3.96
Stack ubuntu 3.38 3.56 3.78
Googlemap CT 3.10 3.30 4.06
Amazon movies 3.78 3.79 4.11
Yelp 3.80 4.23 4.33
GDGB Sephora 4.69 4.69 4.77
Dianping 4.32 4.34 4.71
WikiRevision 4.74 4.79 4.96
WikiLife 4.44 4.46 4.59
IMDB 3.91 4.02 4.44
WeiboTech 4.84 4.88 4.97
WeiboDaily 4.80 4.92 4.99
Cora 3.98 4.10 4.52
Table 13: The results on the graph embedding metric under TDGG with GPT as the LLM backbone.
Benchmark Dataset w/o M. w/ M. w/ M. & R.
DTGB Enron 0.201 0.259 0.400
GDELT 0.255 0.348 0.393
ICEWS1819 0.198 0.289 0.383
Stack elec 0.235 0.319 0.291
Stack ubuntu 0.223 0.327 0.330
Googlemap CT 0.271 0.302 0.462
Amazon movies 0.303 0.359 0.427
Yelp 0.279 0.396 0.411
GDGB Sephora 0.588 0.671 0.654
Dianping 0.351 0.368 0.369
WikiRevision 0.438 0.422 0.454
WikiLife 0.459 0.453 0.463
IMDB 0.432 0.420 0.435
WeiboTech 0.327 0.488 0.698
WeiboDaily 0.681 0.689 0.694
Cora 0.511 0.446 0.465
Table 14: The results of graph structural quality under IDGG with GPT as the LLM backbone.
Benchmark Dataset Degree MMD $\downarrow$ Spectra MMD $\downarrow$ $D_{k}$ $\alpha$ Power-law Validity
DTGB Enron 0.412 0.378 0.132 1.742 $\times$
GDELT 0.489 0.456 0.167 1.891 $\times$
ICEWS1819 0.503 0.412 0.184 1.623 $\times$
Stack elec 0.537 0.459 0.181 1.654 $\times$
Stack ubuntu 0.521 0.488 0.195 1.577 $\times$
Googlemap CT 0.398 0.342 0.112 1.803 $\times$
Amazon movies 0.476 0.401 0.143 1.668 $\times$
Yelp 0.355 0.317 0.078 1.921 $\times$
GDGB Sephora 0.454 0.189 0.147 2.057
Dianping 0.300 0.453 0.069 2.430
WikiRevision 0.267 0.229 0.045 2.050
WikiLife 0.083 0.208 0.099 2.204
IMDB 0.321 0.423 0.252 1.746 $\times$
WeiboTech 0.268 0.201 0.064 1.867 $\times$
WeiboDaily 0.268 0.439 0.156 1.732 $\times$
Cora 0.136 0.244 0.091 2.250
Table 15: The results on average textual quality scores under IDGG with GPT as the LLM backbone. M. and R. denote node memory and reflection mechanism, respectively.
Benchmark Dataset w/o M. w/ M. w/ M. & R.
DTGB Enron 3.51 3.89 3.93
GDELT 3.62 3.82 4.02
ICEWS1819 3.59 3.90 4.25
Stack elec 3.23 3.46 3.64
Stack ubuntu 3.19 3.47 3.73
Googlemap CT 3.22 3.42 3.87
Amazon movies 3.63 3.75 4.22
Yelp 3.53 4.03 4.16
GDGB Sephora 4.58 4.77 4.87
Dianping 4.56 4.71 4.86
WikiRevision 4.39 4.54 4.71
WikiLife 4.28 4.39 4.47
IMDB 4.19 4.29 4.49
WeiboTech 4.60 4.71 4.93
WeiboDaily 4.74 4.83 4.99
Cora 4.27 4.36 4.57
Table 16: The results on the graph embedding metric under IDGG with GPT as the LLM backbone.
Benchmark Dataset w/o M. w/ M. w/ M. & R.
DTGB Enron 0.221 0.278 0.387
GDELT 0.242 0.359 0.388
ICEWS1819 0.234 0.297 0.361
Stack elec 0.243 0.289 0.321
Stack ubuntu 0.239 0.319 0.309
Googlemap CT 0.289 0.333 0.330
Amazon movies 0.294 0.338 0.396
Yelp 0.281 0.369 0.417
GDGB Sephora 0.601 0.603 0.628
Dianping 0.521 0.527 0.505
WikiRevision 0.589 0.614 0.629
WikiLife 0.510 0.534 0.531
IMDB 0.521 0.535 0.563
WeiboTech 0.501 0.514 0.536
WeiboDaily 0.459 0.481 0.502
Cora 0.541 0.552 0.572
Summary.
Collectively, the experimental results demonstrate that GDGB significantly outperforms DTGB across multiple dimensions of generative evaluation. The superior performance on structural metrics, textual quality, and embedding similarity highlights GDGB’s enhanced semantic richness and compatibility with generative modeling frameworks. In contrast, the limited and often noisy textual content in DTGB hinders its effectiveness in supporting high-fidelity DyTAG generation. Therefore, GDGB is not merely an extension of DTGB with longer text, but a purpose-built benchmark for generative DyTAG tasks, where high-quality textual attributes are a core design principle. These findings establish GDGB as a more suitable foundation for advancing research in generative modeling of dynamic text-attributed graphs.
Appendix D Details of Tasks and Metrics
D.1 Details of GAG-General
GAG-General is a generalized agent-based framework designed for both TDGG and IDGG within the DyTAG paradigm. It extends the existing GAG framework (Ji et al., 2025) through key architectural and procedural generalizations, enabling broader applicability and domain adaptability. The full procedure is specified in Algorithms˜1 and 2; here we detail its design and core distinctions from GAG.
The framework proceeds in three stages. First, generalized node formulation initializes node agents, each optionally equipped with Node Memory and a Reflection Mechanism to maintain and summarize interaction history. Unlike GAG, this stage supports both bipartite and non-bipartite graph structures, allowing modeling of diverse network types. Second, during abstracted interaction simulation, agents generate interactions (or new nodes in IDGG) over sequential rounds. Guided by an LLM backbone and informed by memory and reflection, agents produce textual content such as relations and attributes. Crucially, this workflow is abstracted from domain-specific rules, enabling seamless transfer across datasets without customization—a significant departure from GAG’s rigid, task-specific designs. Finally, in the graph projection and evaluation phase, all generated elements are compiled into a dynamic graph and evaluated using our multi-dimensional metrics.
Hence, the key generalizations over GAG lie in three aspects. 1) GAG-General removes the bipartite constraint, supporting arbitrary graph topologies. 2) It replaces GAG’s domain-specific simulation logic with a unified, abstracted process, enhancing reusability across domains. 3) While GAG focuses on a single generation mode, GAG-General explicitly supports both TDGG and IDGG, capturing distinct dynamics of temporal evolution and structural growth.
In sum, GAG-General serves as a demonstration framework that validates the feasibility of generative DyTAG modeling. Its generalizations extend beyond the original GAG’s limitations, offering a more flexible foundation for future work. While our primary contribution is the GDGB benchmark—encompassing datasets, tasks, and metrics—GAG-General illustrates how such benchmarks can drive the development of more generalizable generative frameworks.
D.2 Details of TDGG and IDGG on GAG-General